OpenStudy (anonymous):

2/3x+4=2x. Solve for x. Someone please explain this to me, im doing something wrong.

hero (hero):

Subtract (2x)/3 from both sides to get \(\large4 = 2x - \frac{2x}{3}\) Next multiply 2x by 3/3 to get: \(\large 4 = \frac{6x}{3} - \frac{2x}{3}\) Combine fractions on the right side to get \(\large 4 = \frac{6x - 2x}{3}\) Subtract like terms in the denominator to get: \(4 = \large \frac{4x}{3}\) Multiply both sides by 3 to get \(4 \dot\ 3 = 4x\) \(12 = 4x\) Divide both sides by 4 to get: \(\large\frac{12}{4} = x\) Finish it!

OpenStudy (anonymous):

Okay, so its 3.. correct?

hero (hero):

That was easy, right?

OpenStudy (anonymous):

No.. Fractions confuse me alot!

hero (hero):

Well, just remember that \(\frac{2}{3} x = \frac{2x}{3}\) It's a little easier to work with when you put the x in the numerator.

OpenStudy (anonymous):

Alright thanks. (: